# CUAAudit: Meta-Evaluation of Vision-Language Models as Auditors of Autonomous Computer-Use Agents

### 11 Mar 2026 v1

@@ Abstract.
copyright: noneccs: Human-centered computing Human computer interaction (HCI)$c$s: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologies Computer visionccs: Computing methodologies Machine learningfootnotetext: This work has been accepted to appear at the AAAI 2026 Workshop on Trust and Control in Agentic AI (TrustAgent).
@@ 3.2. Benchmarks
3.3. Calibration and Confidence Assessment
@@
where $N$ is the total number of evaluated tasks.
@@ 3.4. Inter-Model Agreement
where $p_{o}$ denotes the observed agreement rate between two auditors and $p_{e}$ denotes the expected agreement under independence. Values of $\kappa$ range from 1-1 to $1$, with higher values indicating stronger agreement and $\kappa=0$ corresponding to chance-level agreement.
@@ 4. Results
In this section, we present evaluation of $5$ VLMs as an auditors of CUA across three operating systems (macOS, Windows and Linux). Our analysis focuses on three complementary aspects: (i) accuracy of task completion assessment, (ii) calibration of confidence estimates, and (iii) inter-model agreement.

### 12 Mar 2026 v2

@@ Abstract.
copyright: noneccs: Human-centered computing Human computer interaction (HCI)$c$s: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologies Computer visionccs: Computing methodologies Machine learningfootnotetext: This work has been accepted to appear at the HEAL @ CHI 2026 Worshop on Human-centered Evaluation and Auditing of Language Models.
@@ 3.2. Benchmarks
$3.$ Calibration and Confidence Assessment
@@
where NN is the total number of evaluated tasks.
@@ 3.4. Inter-Model Agreement
where $p_{o}$ denotes the observed agreement rate between two auditors and $p_{e}$ denotes the expected agreement under independence. Values of $\kappa$ range from 1-1 to 11, with higher values indicating stronger agreement and $\kappa=0$ corresponding to chance-level agreement.
@@ 4. Results
In this section, we present evaluation of 55 VLMs as an auditors of CUA across three operating systems (macOS, Windows and Linux). Our analysis focuses on three complementary aspects: (i) accuracy of task completion assessment, (ii) calibration of confidence estimates, and (iii) inter-model agreement.